44. If the probability of winning a game is 0.3, what is the probability of losing it?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Answer: Explaination: Probability of winning a game = 0.3 ∴ Probability of losing the game = 1 – Probability of winning the game = 1 – 0.3 = 0.7
Mahatma Patel
2 years ago
Probability of winning a game  = 0.3
Probability of losing a game = 1 - Probability of winning a game = 1- 0.3 = 0.7